### v3

# Causal Inference for a Hidden Treatment

Abstract
In many empirical settings, directly observing a treatment variable may be infeasible although an error-prone surrogate measurement of the latter will often be available. Causal inference based solely on the surrogate measurement is particularly challenging without validation data. We propose a method that obviates the need for validation data by carefully incorporating the surrogate measurement with a proxy of the hidden treatment to obtain nonparametric identification of several causal effects of interest, including the population average treatment effect, the effect of treatment on the treated, quantile treatment effects, and causal effects under marginal structural models. For inference, we provide general semiparametric theory for causal effects identified using our approach and derive a large class of semiparametric efficient estimators with an appealing multiple robustness property. A significant obstacle to our approach is the estimation of nuisance functions which involve the hidden treatment therefore preventing the direct use of standard machine learning algorithms, which we resolve by introducing a novel semiparametric EM algorithm. We examine the finite-sample performance of our method using simulations and an application which aims to estimate the causal effect of Alzheimer’s disease on hippocampal volume using data from the Alzheimer’s Disease Neuroimaging Initiative.
Keywords: EM algorithm, Misclassification error, Semiparametric theory
1 Introduction
In many practical settings where one may wish to evaluate the causal effect of a hypothetical intervention from observational data, it may be challenging or impossible to directly observe the treatment of interest without error. This may arise either because of mis-classification error in ascertaining the treatment status, or because the treatment is inherently latent and cannot conceivably be directly measured. Examples of the first kind abound in health and social sciences; for instance, self-reported dietary intake in nutrition epidemiology is notoriously error-prone, making it challenging to quantify the extent to which a person meets or exceeds recommended daily dietary intake of a given nutrient, a potential intervention of interest. An example of the second kind arises in evaluating the potential causal impact of early onset dementia on subsequent cognitive decline, as no gold standard currently exists for reaching a definitive diagnosis of early onset dementia (Porsteinsson et al.,, 2021). More broadly, unvalidated self-reported treatment may not reflect true treatment status. Hu and Lewbel, (2012) illustrate this by highlighting potential discrepancies between an individual’s actual college attendance and their claimed education level.
The study of measurement error models is longstanding, driven by its broad relevance. A significant portion of this research has focused on continuous treatments, where classical measurement error is routinely assumed. Various strategies have been developed to address measurement error, and Schennach, (2016) provides a comprehensive review of existing measurement error literature. Compared to the well-explored domain of mismeasured continuous treatments, the study of misclassified binary treatments within a causal context has been far less extensive (Bollinger,, 1996). This may stem from the inherent nonclassical nature of misclassification error of a binary treatment, as such error cannot generally be additive, independent of the hidden treatment value, and mean zero simultaneously, other than perhaps in certain degenerate cases (see §LABEL:supp-sec:classical_error of the Supplementary Materials for details). Mahajan, (2006) considers the identification and estimation of nonparametric and semiparametric conditional mean regression models with a misclassified binary regressor, and leverages an instrumental variable to identify and estimate the model. Lewbel, (2007) likewise uses a three-valued intrumental variable to identify and estimate the conditional average treatment effect (CATE) of a misclassified binary treatment. Hu, (2008) extends the framework to a general categorical treatment subject to misclassification via a so-called matrix diagonalization technique to solve a nonlinear system of equations. Allman et al., (2009) propose a comprehensive solution for identifying the distribution of an unobserved discrete variable through multiple measurements, based on array decomposition results. Validation data is another commonly used tool, where both error-prone and error-free variables are observed on a subsample (Chen et al.,, 2005; Sepanski and Carroll,, 1993).
In the absence of auxiliary data, Chen et al., (2009) establish point identification for a nonparametric location-shift model using characteristic functions. They propose sieve maximum likelihood estimators for certain smooth functionals and provide corresponding semiparametric efficiency theory (Shen,, 1997). They also provide a variational characterization of the efficiency bound for their functionals which may be challenging to use in practice for uncertainty quantification due to the lack of a closed-form expression. Another research strand focuses on partial identification rather than point identification. Imai and Yamamoto, (2010) seek to determine bounds for the average treatment effect (ATE) when a binary treatment is measured with differential error, which allows the error to depend on the outcome. DiTraglia and Garcia-Jimeno, (2019) extends the model in Mahajan, (2006) to accommodate endogeneity of the treatment, establishing conditions for partial identification and point identification.
In this paper, we propose a general approach for the identification and semiparametric estimation of a large class of causal effects of a misclassified binary treatment, even in the absence of validation data. However, in addition to the error-prone treatment measurement, referred to as surrogate treatment, one has access to a proxy variable that provides an additional exogenous source of variation associated with the hidden treatment. The main distinction between a surrogate and a proxy lies in their relative level of relevance to the hidden treatment. The degree of misclassification error in the surrogate is known to admit a bound that ensures it outperforms a random guess (although the nature of the bound may be unknown). In contrast, misclassification error in the proxy needs not be similarly bounded. Valid surrogate measurements abound in practice, often in the form of self-report treatment as in the nutrition epidemiology example given above. Proxies are also routinely available; for example, a cognitive function test may serve as a proxy for dementia onset if the two have a non-null association, a less stringent requirement than for a surrogate.
Throughout the paper, we use the ATE as a running example of our proposed approach; while other causal parameters and related theoretical extensions are briefly discussed in the main body of the paper, a detailed discussion is relegated to the Supplementary Materials. Our method for identifying ATE draws on the matrix diagonalization approach detailed by Hu, (2008), which we apply to a broad class of causal parameters. While the identification strategy has a longstanding precedent, the estimation methods proposed by Mahajan, (2006), Lewbel, (2007), and Hu, (2008) have several limitations. Mainly, Lewbel, (2007) adopts a Generalized Method of Moments (GMM) estimator for discrete covariates and employs a local-GMM approach for continuous covariates at a fixed point, limiting the practical utility of their approach. The plug-in estimators proposed by Mahajan, (2006) and Hu, (2008) both follow directly from the identification procedure, making them particularly vulnerable to bias in the estimated law for the observed data distribution. This sensitivity contributes to their instability, as shown in cases where the estimated probabilities involving the hidden treatment may not always be real numbers.
In contrast to the estimators mentioned above, we adopt a modern semiparametric theory framework (Bickel et al.,, 1993; Newey,, 1990; van der Vaart,, 1991) to develop a general method for constructing regular and asymptotically linear (RAL) estimators of various causal parameters of interest. Our results rely upon a novel characterization of a formal mapping between the Hilbert space of all influence functions for the causal functional of interest in the idealized oracle model where misclassification error is absent, and the corresponding space of all influence functions in the observed data model in which the treatment is hidden. This characterization delivers a large class of new multiply robust semiparametric estimators for a rich class of causal parameters in the hidden treatment model. Multiple robustness ensures that our estimators remain consistent even if some but not all required nuisance functions are estimated consistently. This implies that our estimators admit a mixed-bias property (Robins et al.,, 2008; Rotnitzky et al.,, 2021; Ghassami et al.,, 2022) which guarantees root-n consistency and asymptotic normality even if some nuisance functions are estimated nonparametrically at relatively slow rates, provided that other nuisance functions can be estimated at sufficiently fast rates. A unique challenge we face for inference is that nuisance functions involve the hidden treatment, requiring careful estimation. We address this by adopting a novel semiparametric implementation of the EM algorithm inspired by Benaglia et al., (2009). The approach appropriately targets the nuisance parameters without compromising the local robustness property of our treatment effect estimator. Importantly, our approach avoids the need for a consistent estimator of the entire observed data law and bypasses the inverse-problem challenges faced by Mahajan, (2006) and Hu, (2008), thereby ensuring the possibility for multiple robustness which they lack. As we establish, our estimator is semiparametric efficient in the important case of a binary outcome. For more general categorical outcomes, we derive a large class of semiparametric estimators and characterize the efficiency bound in the hidden treatment model. Furthermore, we extend our approach to a rich class of causal functionals, providing multiply robust estimators for both discrete and continuous outcomes, including an approximate semiparametric efficient estimator in the continuous case.
To summarize our key contributions:
1. We unify existing point identification techniques for a large class of causal parameters in hidden treatment models where both proxy and surrogate measurements are available.
2. We present semiparametric theory for several common causal functionals under the hidden treatment model by developing a formal mapping from the oracle tangent space where the hidden treatment is observed, to the observed data tangent space. This mapping enables a characterization of the set of influence functions for the causal parameters of interest.
3. We characterize the semiparametric efficiency bound for the hidden treatment model based on a closed-form expression for the efficient influence function for any causal parameter.
4. We propose a class of multiply robust estimators for functionals in the hidden treatment model which require a novel semiparametric EM algorithm for estimating nuisance functions involving the hidden treatment. Our estimators are discussed in detail in the special case of the ATE, and evaluated via empirical experiments and an application.
The rest of the paper is organized as follows. §2 reviews some common causal parameters of interest. In §3, we introduce the hidden treatment model and present key identification results. §4 develops the semiparametric theory for the ATE, and §5 details the construction of a semiparametric estimator along with its formal asymptotic theory. §6 extends these results to other causal functionals. We illustrate the effectiveness of the method via a simulation study in §7 and an application on the impact of Alzheimer’s disease on hippocampal volume with imperfect diagnosis in §8. The paper concludes with a discussion in §9.
2 Examples of causal parameters
Let $A^{*}$ denote the hidden binary treatment, $Y$ be a binary, count, or continuous outcome, and $X$ be a vector of observed covariates. The potential outcome $Y^{a^{*}}$ represents a unit’s outcome had, possibly contrary to fact, the treatment $A^{*}$ been set to $a^{*}$. Depending on the nature of the outcome and of the scientific inquiry, a variety of causal effects might be of interest. Below we briefly review several examples of causal parameters of broad interest. Throughout, we make standard positivity, consistency and no unmeasured confounding assumptions about the hidden treatment as formally stated below, allowing the exposition to focus solely on the identification and estimation challenges related to treatment misclassification. For each parameter, we briefly review its identifying functional and corresponding class of influence functions that an oracle with access to $A^{*}$ might use as basis for inference.
Example 1.
Average treatment effect (ATE). The ATE measures the average additive effect of a treatment on the outcome, $E(Y^{a^{*}=1}-Y^{a^{*}=0})$. It therefore suffices to identify $E(Y^{a^{*}}),a^{*}=0,1$. The ATE can be identified by the oracle using the standard g-formula
[…]
under the following standard causal assumptions (Robins,, 1986).
Assumption 1.
i. (Ignorability) $Y^{a^{*}}\!\perp\!\!\!\perp A^{*}\mid X$ for $a^{*}=0,1$; ii. (Positivity) $10<\textrm{pr}(A^{*}=1\mid X)<$ 1 almost surely; iii. (Consistency) $Y^{A^{*}}=Y$ almost surely.
Furthermore, the efficient influence function (EIF) for $\psi_{a^{*}}$ in a nonparametric model that places no restriction on the joint law of $(A^{*},Y,X)$ is
$\phi_{a^{*}}^{F}=\frac{\mathbb{I}\left(A^{*}=a^{*}\right)}{f\left(A^{\ast}=a^{*}\mid X\right)}\left\{Y-E\left(Y\mid A^{*}=a^{*},X\right)\right\}+E\left(Y\mid A^{*}=a^{*},X\right)-\psi_{a^{*}},$ (2)
where $f(A^{*}=a^{*}\mid X)$ is the propensity score and $E(Y\mid A^{*},X)$ is the outcome regression function (Robins et al.,, 1994). The upperscript $F$ denotes the full data model where $A^{*}$ is observed.
Remark 1.
For categorical $Y$, the parameters of interest would be $E\{\mathbb{I}(Y^{a^{*}}=y)\}$ where $y$ is any category $Y$ can assume. Its EIF can be obtained by replacing $Y$ with $\mathbb{I}(Y=y)$ in (2).
Example 2.
The Effect of Treatment on the Treated (ETT). The ETT is defined as $\psi_{ett}=E\{Y^{a^{*}=1}-Y^{a^{*}=0}\mid A^{*}=1\}$. It is often of interest when evaluating the effect of treatments that require voluntary participation from members of the population. Under Assumption 1, the ETT is identified by $\psi_{ett}=E\{E(Y\mid A^{*}=1,X)-E(Y\mid A^{*}=0,X)\mid A^{*}=1\}$. The EIF for $\psi_{ett}$ as derived in Hahn, (1998) is given by
$\phi^{F}_{ett}$ $=\frac{\mathbb{I}(A^{*}=1)}{\textrm{pr}(A^{*}=1)}\{Y-E(Y\mid A^{*}=1,X)\}-\frac{\mathbb{I}(A^{*}=0)}{\textrm{pr}(A^{*}=1)}\frac{\textrm{pr}(A^{*}=1\mid X)}{1-\textrm{pr}(A^{*}=1\mid X)}\{Y-E(Y\mid A^{*}=0,X)\}$
$\quad+\frac{E(Y\mid A^{*}=1,X)-E(Y\mid A^{*}=0,X)-\psi_{ett}}{\textrm{pr}(A^{*}=1)}\textrm{pr}(A^{*}=1\mid X).$ (3)
Example 3.
Quantile treatment effect (QTE). Quantile treatment effects examine the impact of a treatment on a given quantile of the outcome (e.g., the median), rather than focusing on the treatment effect on the average outcome. Suppose we are interested in the $\gamma$-quantile of a continuous potential outcome $Y^{a^{*}}$, i.e., the $\theta_{1}^{a^{*}}$ such that $\textrm{pr}(Y^{a^{*}}\leq\theta_{1}^{a^{*}})=\gamma$. Under Assumption 1, the QTE can be identified. The EIF for $\theta_{1}$ is proportional to
$\phi_{a^{*},\gamma}^{F}=\mathbb{I}(A^{*}=a^{*})\frac{\mathbb{I}(Y\leq\theta_{1}^{a^{*}})-\eta_{1}(X;\theta_{1}^{a^{*}})}{\eta_{2}(a^{*};X)}+\eta_{1}(X;\theta_{1}^{a^{*}})-\gamma,$ (4)
where $\eta_{1}(X;\theta_{1}^{a^{*}})=\textrm{pr}(Y\leq\theta_{1}^{a^{*}}\mid X,A^{*}=a^{*})$ and $\eta_{2}(a^{*};X)=\textrm{pr}(A^{*}=a^{*}\mid X)$.
Example 4.
Semi-linear structural regression model. One flexible model for estimating treatment effects while controlling for confounders is $E(Y\mid A^{*},X)=\beta_{x}(X)+\beta_{a^{*}}A^{*}$, where $\beta_{x}(\cdot)$ is an unknown function of $X$. Any influence function for this model has the form
$\phi^{F}_{sl}=h(X)\{Y-\beta_{a^{*}}A^{*}-\beta_{x}(X)\}\{A^{*}-\textrm{pr}(A^{*}=1\mid X)\},$ (5)
where $h(X)$ is any user-specified function with finite variance.
Example 5.
Marginal structural generalized linear model. Consider the model $g\{E(Y^{a^{*}}\mid V)\}=\beta_{0}+\beta_{1}a^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V$ where $g$ is a canonical link function and $V$ is a strict subset of the covariates $X$. Any influence function for this model must be of the form
$\phi^{F}_{msm}$ $=Wh(A^{*},V)\left\{Y-g^{-1}(\beta_{0}+\beta_{1}A^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\}-Wh(A^{*},V)\left\{E(Y\mid A^{*},X)-g^{-1}(\beta_{0}+\beta_{1}A^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\}$
$\quad+\sum_{a^{*}}(-1)^{1-a^{*}}h(a^{*},V)\left\{E(Y\mid A^{*}=a^{*},X)-g^{-1}(\beta_{0}+\beta_{1}a^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\},$ (6)
where $W=1/\textrm{pr}(A^{*}\mid X)$, and $h(A^{*},V)$ is a user-specified function with the same dimension as $(\beta_{0},\beta_{1},\beta_{2}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$.
Because influence functions have zero mean only at the truth, in principle they provide a convenient approach for constructing semiparametric estimators. For instance, in Example 1, by solving the estimating equation $\mathbb{P}_{n}\{\widehat{\phi}_{a^{*}}^{F}(\widehat{\psi}_{a^{*}})\}=0$ based on an estimate of the EIF (2), one recovers the well-known augmented-inverse-probability-weighted estimator (AIPW):
$\widehat{\psi}_{a^{*}}=\mathbb{P}_{n}\left\{\frac{\mathbb{I}\left(A^{*}=a^{*}\right)}{\hat{f}\left(A^{\ast}=a^{*}\mid X\right)}\left\{Y-\hat{E}\left(Y\mid a^{*},X\right)\right\}+\hat{E}\left(Y\mid a^{*},X\right)\right\},$ (7)
where $\hat{f}(A^{\ast}=a^{*}\mid X)$ and $\hat{E}(Y\mid a^{*},X)$ are estimated propensity score and outcome regression function, respectively. This estimator is known to be doubly robust (Scharfstein et al.,, 1999), in that $\widehat{\psi}_{a^{*}}$ is consistent if either $\hat{f}(A^{\ast}=a^{*}\mid X)$ or $\hat{E}(Y\mid a^{*},X)$ is consistent, but not necessarily both. When both are consistent, the estimator achieves the semiparametric efficiency bound for the nonparametric model.
As $A^{*}$ is not directly observed in our setting, the estimator (7) is not feasible, and $\psi_{a^{*}}$ cannot be identified without additional conditions. In the next section, we provide a general approach towards identification correcting for the misclassification error using two auxiliary variables: a surrogate $A$ and a proxy $Z$ of the hidden treatment. Instead of observing $A^{*}$, we assume access to a misclassified measurement $A$. For example, if $A^{*}$ indicates whether a patient is taking a medication, $A$ could be self-reported or extracted from medical records, both prone to errors, such as when a non-compliant patient’s recorded treatment does not match their actual intake. We refer to such an error-prone measurement as a surrogate variable as it attempts to measure the hidden treatment. In contrast, a proxy $Z$ of treatment is only required to be associated with the hidden treatment, although it may not be an explicit measurement of the treatment. These definitions are formalized in §3.
3 Identification of ATE
[…]
In order to identify $\psi_{a^{*}}$, our hidden treatment model imposes Assumptions 2 and 3 below.
Assumption 2 (Relevance).
[…]
Assumption 3 (Conditional independence).
$Y$, $A$, and $Z$ are mutually independent conditional on $(A^{*},X)$. That is, $Y\!\perp\!\!\!\perp A\!\perp\!\!\!\perp Z\mid A^{*},X$.
[…]
Figure 1: (a) $A^{*}$ is the hidden binary exposure, $A$ is a binary surrogate, $Z$ is a binary proxy, $Y$ is the outcome, and $X$ is the confounder. (b) A DAG for the oracle nonparametric model $\mathcal{M^{*}}_{np}$.
[…]
Remark 2.
We implicitly assume that $A\neq A^{*}$ and $Z\neq A^{*}$, a necessary implication of misclassification error. If $A$ equals $A^{*}$, the matrix $P(Y,Z,A=1)$ which will be introduced shortly, would not be of full rank. Therefore, a non-zero determinant of $P(Y,Z,A=1)$ indicates that $A\neq A^{*}$. A similar argument applies to $Z$.
Under Assumptions 2 and 3, the joint distribution of $(A^{*},Y,A,Z,X)$ is determined up to labeling. To illustrate the idea, consider a simplified scenario where $Y$ is binary and $X=\emptyset$. In this situation, the joint distribution of the observed data $(Y,A,Z)$ has $72^{3}-1=$ 7 degrees of freedom, and we have 7 unknown parameters: $E(A^{*}),E(Y\mid A^{*}=a^{*}),E(A\mid A^{*}=a^{*}),E(Z\mid A^{*}=a^{*}),a^{*}=0,1$. Let $P(Y,Z)=(\textrm{pr}(Y=2-i,Z=2-j))_{i,j}$, $P(Y,Z,A=1)=(\textrm{pr}(Y=2-i,Z=2-j,A=1))_{i,j}$, $P(Y\mid A^{*})=(\textrm{pr}(Y=2-i\mid A^{*}=2-j))_{i,j}$, $P(Z\mid A^{*})=(\textrm{pr}(Z=2-i\mid A^{*}=2-j))_{i,j}$, $P_{D}(A^{*})=\text{diag}(\textrm{pr}(A^{*}=1),\textrm{pr}(A^{*}=0))$, $P_{D}(A=1\mid A^{*})=\text{diag}(\textrm{pr}(A=1\mid A^{*}=1),\textrm{pr}(A=1\mid A^{*}=0))$. By Assumption 3, we have $P(Y,Z)=P(Y\mid A^{*})P_{D}(A^{*})P(Z\mid A^{*})^{\mathrm{\scriptscriptstyle T}},P(Y,Z,A=1)=P(Y\mid A^{*})P_{D}(A^{*})P_{D}(A=1\mid A^{*})P(Z\mid A^{*})^{\mathrm{\scriptscriptstyle T}}.$ By Assumption 2, $P(A\mid Z)$ is invertible. Multiplying $P(Y,Z,A=1)$ by $P(A\mid Z)^{-1}$ gives $P(Y,Z,A=1)P(Y,Z)^{-1}=P(Y\mid A^{*})P_{D}(A=1\mid A^{*})P(Y\mid A^{*})^{-1},$ which as noted in Hu, (2008) is the canonical form of the eigendecomposition of the matrix on the left-hand side. Since the right-hand side is observed, we can solve for the eigenvalues and eigenvectors, which correspond to the diagonal elements of $P_{D}(A=1\mid A^{*})$ and the columns of $P(Y\mid A^{*})$, respectively. While eigenvectors are identified up to scale, the requirement that columns of $P(Y\mid A^{*})$ sum to 1 fixes it. The remaining challenge lies in correctly matching each eigenvalue with its respective diagonal element in $P_{D}(A=1\mid A^{*})$. To resolve this, we propose the following assumption encoding four alternative ways in which the surrogate $A$ carries sufficient information about $A^{*}$ to resolve the remaining identification gap.
Assumption 4 (Surrogate).
At least one of the following holds: for any value $x$ of $X$,
  1. Condition 1
    . $\textrm{pr}(A=0\mid A^{*}=1,X=x)+\textrm{pr}(A=1\mid A^{*}=0,X=x)<1$
  2. Condition 2
    . $\textrm{pr}(A^{*}=0\mid A=1,X=x)+\textrm{pr}(A^{*}=1\mid A=0,X=x)<1$
  3. Condition 3
    . $\textrm{pr}(A^{*}=1\mid A=1,X=x)>\textrm{pr}(A^{*}=0\mid A=1,X=x)$
  4. Condition 4
    . $\textrm{pr}(A^{*}=0\mid A=0,X=x)>\textrm{pr}(A^{*}=1\mid A=0,X=x)$
Assumption 4 clarifies the different roles of $A$ and $Z$, although they appear symmetric in Figure 1(a). While both the proxy $Z$ and the surrogate $A$ need to satisfy Assumptions 2 and 3, an extra Assumption 4 is imposed on the surrogate. This assumption essentially suggests that $A$, as a surrogate, should provide more information about $A^{*}$ than a random guess. The condition is expected to hold if $A$ is a meaningful but imperfect measurement of $A^{*}$. Condition 1 is equivalent to $\textrm{pr}(A=1\mid A^{*}=0,X=x)<\textrm{pr}(A=1\mid A^{*}=1,X=x)$. Similarly, Condition 2 can be written as $\textrm{pr}(A^{*}=1\mid A=0,X=x)<\textrm{pr}(A^{*}=1\mid A=1,X=x)$. Condition 1 and Condition 2 are effectively of similar but symmetric nature. Condition 3 indicates the precision $\textrm{pr}(A^{*}=1\mid A=1,X=x)>0.5$, while Condition 4 states that the false omission rate $\textrm{pr}(A^{*}=1\mid A=0,X=x)<0.5$. These conditions are reasonable in many scenarios. For example, consider $A^{*}$ as the actual COVID-19 vaccination status of a patient, and $A$ as whether the patient was vaccinated at a specific hospital. If hospital records show a positive vaccination status, it must be that the patient was truly vaccinated, making $\textrm{pr}(A^{*}=1\mid A=1,X=x)=1$, thereby fulfilling Condition 2 (and Condition 1). If one can assume that the misclassification error is balanced across covariates, such that $X\!\perp\!\!\!\perp A\mid A^{*}$, i.e., the surrogate $A$ only depends on $A^{*}$, then one can safely omit $X$ from Assumption 4. We note that Hu, (2008) considers alternatives to Conditions 1-4.
The informal proof given above is formalized in Theorem 1 below in a more general setting where $Y$ can be either binary, a count, or a continuous outcome, and $X$ is possibly vector-valued. For the proof of this theorem, please refer to §LABEL:supp-sec:thm_identification in the supplement.
Theorem 1.
Under Assumptions 1, 2, 3, and 4, the average treatment effect of $A^{*}$ on $Y$, $E(Y^{a^{*}})$, $a^{*}=0,1$, are uniquely identified and therefore $\psi_{a^{*}}=E(Y^{a{{}^{*}}=1}-Y^{a{{}^{*}}=0})$ is uniquely identified whether $Y$ is binary, a count, or a continuous outcome.
A key step in the proof entails establishing point identification of the joint conditional oracle law $\textrm{pr}(A,Z,Y,A^{*}\mid X=x)=\textrm{pr}(A\mid A^{*},X=x)\textrm{pr}(Z\mid A^{*},X=x)\textrm{pr}(Y\mid A^{*},X=x)\textrm{pr}(A^{*}\mid X=x)$, which, in principle implies point identification of any smooth functional of the oracle law. This observation is crucial to establishing identification for more general causal functionals beyond the ATE. The next section builds on this key identification result to construct highly robust and efficient semiparametric estimators of $\psi_{a^{*}}$ under the semiparametric model $\mathcal{M}$ for the observed data, induced by the oracle semiparametric model $\mathcal{M^{*}}$ defined by the conjunction of Assumptions 14.
4 Semiparametric theory
Having established identification for $\psi_{a^{*}}$, our next objective is to characterize the tangent space of the semiparametric model $\mathcal{M}$ and its ortho-complement. By using some key results from modern semiparametric theory, we obtain a set of influence functions for $\psi_{a^{*}}$, which in turn motivates a feasible set of estimating equations. This step is crucial for developing efficient semiparametric estimators for $\psi_{a^{*}}$ and conducting principled inferences under $\mathcal{M}$.
Specifically, our approach starts with a characterization of the tangent space in the oracle full data model, in which $A^{*}$ is fully observed, before transitioning to the observed data model where $A^{*}$ is hidden. This is done by establishing the relevant coarsening operator which maps any oracle score function under $\mathcal{M^{*}}$ to a corresponding score function under $\mathcal{M}$. As we demonstrate, this coarsening operator induces a formal characterization of the tangent space and its ortho-complement in the hidden treatment model. This is finally used to generate a set of influence functions, giving rise to a corresponding rich class of feasible moment equations with appealing robustness and efficiency properties.
Under Assumptions 2 and 3, the full data distribution is $f\left(Y,A,A^{*},Z,X\right)=f\left(Y\mid A^{*},X\right)\allowbreak f\left(A\mid A^{*},X\right)\allowbreak f\left(Z\mid A^{*},X\right)f\left(A^{*}\mid X\right)f\left(X\right),$ with the corresponding tangent space given by
$\Lambda^{F}=\left\{\Lambda_{y}^{F}\oplus\Lambda_{a}^{F}\oplus\Lambda_{z}^{F}\oplus\Lambda_{a^{*}}^{F}\oplus\Lambda_{x}\right\}\cap L_{0}^{2,F},$ (8)
where $\Lambda_{y}^{F}=\left\{G_{y}=g_{y}\left(Y,A^{*},X\right):E\left(G_{y}\mid A^{*},X\right)=0\right\}$, $\Lambda_{a}^{F}=\left\{G_{a}=g_{a}\left(A,A^{*},X\right):E\left(G_{a}\mid A^{*},X\right)=0\right\}$, $\Lambda_{z}^{F}=\left\{G_{z}=g_{z}\left(Z,A^{*},X\right):E\left(G_{z}\mid A^{*},X\right)=0\right\}$, $\Lambda_{a^{*}}^{F}=\left\{G_{a^{*}}=g_{a^{*}}\left(A^{*},X\right):E\left(G_{a^{*}}\mid X\right)=0\right\}$, $\Lambda_{x}=\left\{G_{x}=g_{x}\left(X\right):E\left(G_{x}\right)=0\right\}$. We use $L^{2}_{0}(V)$ to denote the Hilbert space of measurable functions of $V$ with mean zero and finite variance equipped with the covariance inner product, and $L_{0}^{2,F}$ is a shorthand for $L^{2}_{0}(Y,A,A^{*},Z,X)$.
It can be verified that the orthocomplement of the tangent space (8) is represented by
$\begin{split}\Lambda^{F,\bot}=\bigl{\{}K=&k\left(Y,A,A^{*},Z,X\right):E\left(K\mid Y,A^{*},X\right)\\ &=E\left(K\mid A,A^{*},X\right)=E\left(K\mid Z,A^{\ast},X\right)=0\bigr{\}}\cap L_{0}^{2,F},\end{split}$ (9)
which can be equivalently expressed as
$\begin{split}\Lambda^{F,\bot}=\bigl{\{}W=M-&E\left(M\mid Y,A^{*},X\right)-E\left(M\mid A,A^{*},X\right)-E\left(M\mid Z,A^{*},X\right)\\ &+2E\left(M\mid A^{\ast},X\right):M=m\left(Y,A,A^{*},Z,X\right)\in L_{0}^{2,F}\bigr{\}}.\end{split}$ (10)
The derivation of (9) and (10) is deferred to §LABEL:supp-sec:proof_tangentspace_full in the Supplementary Materials. In the following, we will make reference to the oracle nonparametric model $\mathcal{M}_{np}^{*}$ as the collection of oracle laws compatible with the DAG given in Figure 1(b) which places no restriction on the joint distribution of $(A^{*},Y,Z,A,X)$. We then have the following result:
Lemma 1.
The EIF of $\psi_{a^{*}}$ in the oracle hidden treatment model $\mathcal{M^{*}}$ is equal to the EIF in the oracle nonparametric model $\mathcal{M}_{np}^{*}$.
The proof of Lemma 1 is relegated to the Supplementary Materials, §LABEL:supp-sec:eif_lemma.
The EIF for $\psi_{a^{*}}$ in $\mathcal{M^{*}}_{np}$, denoted by $\phi_{a^{*}}^{F,eff}$ is well-established, and is equal to (2). Therefore, the set of influence functions for $\psi_{a^{*}}$ in the full data model $\mathcal{M^{*}}$ is the linear variety $\phi_{a^{*}}^{F,eff}\oplus\Lambda^{F,\bot}$ (Tsiatis,, 2007).
So far, we have characterized the tangent space, the set of influence functions and the EIF for $\psi_{a^{*}}$ in $\mathcal{M^{*}}$. We now turn to the observed data model $\mathcal{M}$. The following result summarizes our characterization of the tangent space of the observed data hidden treatment model $\Lambda$, and its ortho-complement, obtained by applying a certain coarsening operator which maps the tangent space for the full data model, $\Lambda^{F}$, onto the $L^{2}_{0}(Y,A,Z,X)$ space.
Theorem 2.
The tangent space $\Lambda$ for the observed data hidden treatment model is given by $\Lambda=\{\Lambda_{y}+\Lambda_{a}+\Lambda_{z}+\Lambda_{a^{*}}+\Lambda_{x}\}\cap L^{2}_{0}$, where $\Lambda_{y}=E(\Lambda_{y}^{F}\mid Y,A,Z,X)$, $\Lambda_{a}=E(\Lambda_{a}^{F}\mid Y,A,Z,X)$, $\Lambda_{z}=E(\Lambda_{z}^{F}\mid Y,A,Z,X)$, $\Lambda_{a^{*}}=E(\Lambda_{a^{*}}^{F}\mid Y,A,Z,X)$, $\Lambda_{x}=E(\Lambda_{x}^{F}\mid Y,A,Z,X)$. The ortho-complement of the observed data tangent space, $\Lambda^{\bot}$ is characterized by
$\Lambda^{\bot}=\{L=l(Y,A,Z,X):E(L\mid Y,A^{*},X)=E(L\mid A,A^{*},X)=E(L\mid Z,A^{*},X)=0\}\cap L^{2}_{0},$
which is equivalently expressed as
$\begin{split}&\Lambda^{\bot}=\bigl{\{}\{A-E(A\mid A^{*}=0,X)\}\{Z-E(Z\mid A^{*}=1,X)\}h_{1}(Y,X)+\{A-E(A\mid A^{*}=1,X)\}\\ &\{Z-E(Z\mid A^{*}=0,X)\}h_{2}(Y,X):E\{h_{1}(Y,X)\mid A^{*},X\}=E\{h_{2}(Y,X)\mid A^{*},X\}=0\bigr{\}}\cap L^{2}_{0}.\end{split}$ (11)
We illustrate the theorem in key special cases. Thus, suppose $Y$ is a categorical variable with levels $1,\ldots,k$, the functions $h_{1}(Y,X)$ and $h_{2}(Y,X)$ must fulfill the following condition:
$\begin{bmatrix}h_{j}(Y=1,X)&\cdots&h_{j}(Y=k,X)\end{bmatrix}M_{Y\mid A^{*},X}=\begin{bmatrix}0&0\end{bmatrix},\quad j=1,2$ (12)
where $M_{Y\mid A^{*},X}$ is defined as $(v_{1},v_{2})$, with $v_{1}=\left(f(Y=1\mid A^{*}=0,X),\cdots,f(Y=k\mid A^{*}=0,X)\right)^{\mathrm{\scriptscriptstyle T}}$ and $v_{2}=\left(f(Y=1\mid A^{*}=1,X),\cdots,f(Y=k\mid A^{*}=1,X)\right)^{\mathrm{\scriptscriptstyle T}}$. This implies that $h_{1}$ and $h_{2}$ belong to the null space of the $k\times 2$ matrix $M_{Y\mid A^{*},X}$. Under Assumption 2, the null space has a dimension of $k-2$ when $k\geq 2$. When $Y$ is binary, i.e., $k=2$, $M_{Y\mid A^{*},X}$ is full rank, leading to the unique solution of (12): $h_{j}(Y,X)=0,j=1,2$. This is further detailed in Corollary 1.
Corollary 1.
Suppose that $A^{*},A,Z,Y$ are all binary variables, then the observed data hidden treatment model is locally nonparametric in the sense that the tangent space of $\mathcal{M}$ is $L^{2}_{0}$, i.e., $\Lambda=\{L=l(Y,A,Z,X)\in L^{2}_{0}\}$.
For more general $Y$, i.e., categorical or continuous, the following theorem provides a closed-form characterization of $\Lambda$ which facilitates influence function-based inference. In this vein, let $P_{1}=\{h(Y,X):E\{h(Y,X)\mid A^{*},X\}=0\}\cap L^{2}_{0}$ and the space of functions
$P_{2}=\Biggl{\{}v(Y,X)=\frac{v_{1}(Y,X)-E\{v_{1}(Y,X)\mid A^{*}=0,X\}}{E\{v_{1}(Y,X)\mid A^{*}=1,X\}-E\{v_{1}(Y,X)\mid A^{*}=0,X\}}$
$\qquad\qquad+\frac{v_{0}(Y,X)-E\{v_{0}(Y,X)\mid A^{*}=1,X\}}{E\{v_{0}(Y,X)\mid A^{*}=0,X\}-E\{v_{0}(Y,X)\mid A^{*}=1,X\}}-1:v_{1}(Y,X),v_{0}(Y,X)\in L^{2}\Biggr{\}}.$
Theorem 3.
Suppose that $Y$ is bounded, i.e., $|Y|<L$; then we have that $P_{1}=P_{2}$; and
$\Lambda^{\bot}=\bigl{\{}\{A-E(A\mid A^{*}=0,X)\}\{Z-E(Z\mid A^{*}=1,X)\}w_{1}(Y,X)+\{A-E(A\mid A^{*}=1,X)\}$
$\qquad\quad\,\{Z-E(Z\mid A^{*}=0,X)\}w_{2}(Y,X):w_{1},w_{2}\in P_{2}\bigr{\}}.$ (13)
Such equivalence provides a closed-form characterization of $P_{1}$. Based on this, we obtain a closed-form characterization of $\Lambda^{\bot}$. See §LABEL:supp-sec:orthocomplement in the Supplementary Materials for the proof.
Having characterized the observed data tangent space, we note that the set of all observed data influence functions for the functional $\psi_{a^{*}}$ in the hidden treatment model can be written as $\phi+\Lambda^{\bot}$ for any influence function $\phi$. Therefore, it suffices to identify a single influence function $\phi$. To accomplish this, we establish a formal relationship between influence functions in the observed and oracle data models. As we have characterized the influence functions for the oracle model, we may leverage this relationship to derive the corresponding set of influence functions in the observed hidden treatment model. The following result provides such a characterization of observed data influence functions of $\psi_{a^{*}}$.
Lemma 2.
Suppose that $Q$ is an influence function of $\psi_{a^{*}}$ in $\mathcal{M}$, then, there must exist an oracle influence function of $\psi_{a^{*}}$ in $\mathcal{M^{*}}$ denoted by $\text{IF}^{F}$ such that:
$\begin{split}E\left(Q\mid Y,A^{*},X\right)&=E\left(\text{IF}^{F}\mid Y,A^{*},X\right),\\ E\left(Q\mid A,A^{*},X\right)&=E\left(\text{IF}^{F}\mid A,A^{*},X\right),\\ E\left(Q\mid Z,A^{*},X\right)&=E\left(\text{IF}^{F}\mid Z,A^{*},X\right).\end{split}$ (14)
Once we have identified such a $Q$, the set of influence functions for $\psi_{a^{*}}$ in the observed data model is given by $Q+\Lambda^{\bot}$. Theorem 4 provides such an influence function for the observed data model.
Theorem 4.
The following equation provides a closed-form expression of an influence function for $\psi_{a^{*}}$ in the hidden treatment model $\mathcal{M}$ which applies whether $Y$ is binary, a count or a continuous variable:
$\phi_{a^{*}}(Y,A,Z,X)=$ $\frac{A-E(A\mid 1-a^{*},X)}{E(A\mid a^{*},X)-E(A\mid 1-a^{*},X)}\frac{Z-E(Z\mid 1-a^{*},X)}{E(Z\mid a^{*},X)-E(Z\mid 1-a^{*},X)}\frac{1}{f(a^{*}\mid X)}$
$\{Y-E(Y\mid a^{*},X)\}+E(Y\mid a^{*},X)-\psi_{a^{*}},$ (15)
with $Y$ replaced by $\mathbb{I}(Y=y)$ for a categorical $Y$ where $\psi_{a^{*}}=E\{\mathbb{I}(Y^{a^{*}}=y)\}$. Equation (4) admits the triple robustness property: $E(\phi_{a^{*}})=0$ if one of the following holds:
1. $f(A\mid A^{*},X)$ and $f(Y\mid A^{*},X)$ are correctly specified.
2. $f(Z\mid A^{*},X)$ and $f(Y\mid A^{*},X)$ are correctly specified.
3. $f(A\mid A^{*},X)$, $f(Z\mid A^{*},X)$ and $f(A^{*}\mid X)$ are correctly specified.
The EIF $\phi_{a^{*}}^{eff}$ is given by the orthogonal projection of $\phi_{a^{*}}$ onto the observed data tangent space of the hidden treatment model, i.e., $\phi_{a^{*}}^{eff}=\Pi(\phi_{a^{*}}\mid\Lambda)$ where $\Pi$ is the orthogonal projection operator. Therefore,
$\phi_{a^{*}}^{eff}$ $=\phi_{a^{*}}(Y,A,Z,X)-\{A-E(A\mid A^{*}=0,X)\}\{Z-E(Z\mid A^{*}=1,X)\}w^{opt}_{1}(Y,X)$
$\qquad-\{A-E(A\mid A^{*}=1,X)\}\{Z-E(Z\mid A^{*}=0,X)\}w^{opt}_{2}(Y,X).$
In the case of binary $Y$, $w^{opt}_{1}=w^{opt}_{2}=0$, and thus (4) is the EIF for $\psi_{a^{*}}$ in the hidden treatment model $\mathcal{M}$. A closed-form expression for $\{w^{opt}_{j}:j=1,2\}$ in the more general categorical outcome setting is given in the Supplemental Materials §LABEL:supp-sec:closed_form_w. As a closed-form expression for $\{w^{opt}_{j}:j=1,2\}$ and thus for the EIF is not generally available for continuous $Y$, we give in the same Supplemental Materials Section a simple approach for obtaining an approximate EIF based on a basis function approximation of $\{w^{opt}_{j}:j=1,2\}$.
The theorem reveals an insightful relationship between an oracle influence function and the corresponding observed data influence function. Specifically, comparing (4) with the full data influence function (2), we observe that the former is obtained by substituting the indicator function $\mathbb{I}_{a^{*}}(A^{*})$ in the latter with:
$M(A,Z,X)=\frac{A-E(A\mid 1-a^{*},X)}{E(A\mid a^{*},X)-E(A\mid 1-a^{*},X)}\frac{Z-E(Z\mid 1-a^{*},X)}{E(Z\mid a^{*},X)-E(Z\mid 1-a^{*},X)}.$ (16)
Interestingly, (16) is a conditionally unbiased estimator of $\mathbb{I}_{a^{*}}(A^{*})$, satisfying $E\{M(A,Z,X)\mid A^{*},X\}=\mathbb{I}_{a^{*}}(A^{*})$. See §LABEL:supp-sec:relationship_if for the derivation. This provides an elegant link between (2) and (4). In (2), the unobserved $A^{*}$ only appears within the indicator function $\mathbb{I}_{a^{*}}(A^{*})$, while in (4), $\mathbb{I}_{a^{*}}(A^{*})$ is replaced by an expression of observed variables (16), whose expectaction conditional on $(A^{*},X)$ equals $\mathbb{I}_{a^{*}}(A^{*})$. This observation suggests a general approach for deriving influence functions that can be extended to other causal functionals of $(A^{*},X,Y)$ discussed in §2 under the hidden treatment model. See §6 for a detailed exposition.
5 Semiparametric estimation and inference
In this section, we focus on developing nonparametric estimators for $\psi_{a^{*}}$, based on the influence function (4). A standard approach involves the estimation of nuisance parameters, which are then substituted in the influence function to obtain a substitution estimator:
$\begin{split}\widehat{\psi}_{a^{*}}=\mathbb{P}_{n}\Biggl{\{}\frac{A-\widehat{E}(A\mid 1-a^{*},X)}{\widehat{E}(A\mid a^{*},X)-\widehat{E}(A\mid 1-a^{*},X)}\frac{Z-\widehat{E}(Z\mid 1-a^{*},X)}{\widehat{E}(Z\mid a^{*},X)-\widehat{E}(Z\mid 1-a^{*},X)}\\ \frac{1}{\widehat{f}(a^{*}\mid X)}\{Y-\widehat{E}(Y\mid a^{*},X)\}+\widehat{E}(Y\mid a^{*},X)\Biggr{\}},\end{split}$ (17)
where $\mathbb{P}_{n}$ is the empirical mean operator.
Typically, nuisance functions are estimated in an initial step from the observed data. However, because the nuisance functions in (17) involve the unobserved $A^{*}$, their estimation is challenging, especially when using nonparametric estimators. One approach is to first estimate the observed data distribution and then compute nuisance functions via a standard plug-in principle. For a detailed discussion, see §LABEL:supp-sec:me_estimator of the Supplementary Materials. A key challenge of this approach is its heavy dependence on the accuracy of estimating the observed data distribution $\textrm{pr}(Y,A,Z\mid X)$, as the nuisance functions are solutions to polynomial equations that are smooth in the observed data law. This dependence renders the estimated nuisance functions highly sensitive to the uncertainty in $\widehat{\textrm{pr}}(Y,A,Z\mid X)$, complicating efforts to obtain estimates of the nuisance functions that are guaranteed to remain in the model. To overcome this, we propose a semiparametric EM algorithm that circumvents the need for an initial estimator of the observed data law. Our algorithm is inspired by the nonparametric EM algorithm in Benaglia et al., (2009). See Algorithm 1 for implementation details.
The initial values $\varphi^{(0)}_{i},i=1,\ldots,n$, serve as starting point and can be randomly assigned within a sensible range. The inputs $\widehat{\textrm{pr}}(Y\mid X)$, $\widehat{\textrm{pr}}(A\mid X)$, $\widehat{\textrm{pr}}(Z\mid X)$ are the estimated conditional probabilities of $Y,A,Z$ given $X$, respectively. They can be obtained by user-specified nonparametric methods such as kernel regression, local polynomial regression, etc. Throughout our simulation and real data example, we employ kernel regression with a Gaussian kernel, and the bandwidth $h$ is calculated from the Silverman’s rule-of-thumb formula $h=1.06\text{sd}(X)n^{-1/(4+d)}$ (Silverman,, 1986), with $d$ being the dimension of the regressor.
In practice, estimation processes often incorporate cross-fitting to prevent over-fitting (Chernozhukov et al.,, 2018; Schick,, 1986) by using separate parts of the dataset for estimating nuisance functions and calculating the empirical mean in (17). We employ a modified cross-fitting approach that slightly deviates from the standard procedure, as the training and evaluation sets are iterated simultaneously. Specifically, during each M-step, models $\widehat{E}(p^{(t)}\mid X),\widehat{E}(p^{(t)}\mid Y,X),\widehat{E}(p^{(t)}\mid A,X),\widehat{E}(p^{(t)}\mid Z,X)$ are fitted using the training set. These models are then applied to both the training and evaluation set to obtain $\widetilde{\textrm{pr}}(A^{*}=1\mid X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid Y_{i},X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid A_{i},X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid Z_{i},X_{i})$ for $i=1,\ldots,n$. See Algorithm LABEL:supp-alg:EM_cf in the Supplementary Materials for details. Notably, the models evaluating the nuisances are independent of the evaluation data, making it a valid form of cross-fitting.
  • Input $\varphi^{(0)}_{i}=(\textrm{pr}^{(0)}(A^{*}\mid X),\textrm{pr}^{(0)}(Y\mid A^{*},X),\textrm{pr}^{(0)}(A\mid A^{*},X),\textrm{pr}^{(0)}(Z\mid A^{*},X))_{i},i=1,\ldots,n$ and $(\widehat{\textrm{pr}}(Y\mid X)$, $\widehat{\textrm{pr}}(A\mid X)$, $\widehat{\textrm{pr}}(Z\mid X))_{i},i=1,\ldots,n$.
For $t=1,2,\ldots$, repeat 1 E-step: Calculate the “posterior” probabilities conditional on the data and $(\varphi^{(t)}_{1},\ldots,\varphi^{(t)}_{n})$.
For $i=1,\ldots,n$,
$p^{(t)}_{i}:=$ $\textrm{pr}^{(t)}(A^{*}=1\mid Y_{i},A_{i},Z_{i},X_{i})$
[…]
  • 2
    M-step: Estimate
    $\widetilde{\textrm{pr}}(A^{*}=1\mid X=x)$ $=\widehat{E}(p^{(t)}\mid X=x),$
    $\widetilde{\textrm{pr}}(A^{*}=1\mid Y=y,X=x)$ $=\widehat{E}(p^{(t)}\mid Y=y,X=x),$
    $\widetilde{\textrm{pr}}(A^{*}=1\mid A=a,X=x)$ $=\widehat{E}(p^{(t)}\mid A=a,X=x),$
    $\widetilde{\textrm{pr}}(A^{*}=1\mid Z=z,X=x)$ $=\widehat{E}(p^{(t)}\mid Z=z,X=x).$
  • 3
    Probability update step: For $a^{*}=0,1$, set
    $\textrm{pr}^{(t+1)}(A^{*}=a^{*}\mid X_{i})$ $=\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i}),$
    $\textrm{pr}^{(t+1)}(Y_{i}\mid A^{*}=a^{*},X_{i})$ $=\frac{\widehat{\textrm{pr}}(Y_{i}\mid X_{i})\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid Y_{i},X_{i})}{\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i})},$
    $\textrm{pr}^{(t+1)}(A_{i}\mid A^{*}=a^{*},X_{i})$ $=\frac{\widehat{\textrm{pr}}(A_{i}\mid X_{i})\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid A_{i},X_{i})}{\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i})},$
    $\textrm{pr}^{(t+1)}(Z_{i}\mid A^{*}=a^{*},X_{i})$ $=\frac{\widehat{\textrm{pr}}(Z_{i}\mid X_{i})\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid Z_{i},X_{i})}{\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i})},$
    until iteration $T=\min\{t:\max_{i\in[n]}|p^{(t)}_{i}-p^{(t-1)}_{i}|<\delta\}$, where $\delta$ is a user-defined threshold.
    • Output $\widehat{f}(A^{*}=a^{*}\mid X_{i})=\textrm{pr}^{(T)}(A^{*}=a^{*}\mid X_{i}),$ $\widehat{E}(A\mid A^{*}=a^{*},X_{i})=\textrm{pr}^{(T)}(A=1\mid A^{*}=a^{*},X_{i}),$ $\widehat{E}(Z\mid A^{*}=a^{*},X_{i})=\textrm{pr}^{(T)}(Z=1\mid A^{*}=a^{*},X_{i}),$
      $\widehat{E}(Y\mid A^{*}=a^{*},X_{i})=\begin{cases}\textrm{pr}^{(T)}(Y=1\mid A^{*}=a^{*},X_{i}),&\text{if $Y$ is binary}\\ n^{-1}\sum^{n}_{\ell=1}Y_{\ell}\textrm{pr}^{(T)}(Y=Y_{\ell}\mid A^{*}=a^{*},X_{i}),&\text{if $Y$ is continuous}\end{cases}$
      for $a^{*}=0,1$ and $i=1,\ldots,n$.
  • Algorithm 1 A semiparametric EM algorithm to estimate nuisance parameters in (17)
    The M-step can be implemented in different ways. For example, to estimate $\textrm{pr}(A^{*}=1\mid A=a,X=x)$, one can either use the subset $\{(p^{(t)}_{i},x_{i}):A_{i}=a\}$ to learn $\widetilde{\textrm{pr}}(A^{*}=1\mid A=a,X=x)$, or use the entire dataset $\{(p^{(t)}_{i},a_{i},x_{i}):i=1,\ldots,n\}$ for the same purpose. This concept aligns with the ideas behind the T-learner and S-learner (Künzel et al.,, 2019). However, when $Y$ is continuous, then it is preferable to use the dataset $\{(p^{(t)}_{i},y_{i},x_{i}):i=1,\ldots,n\}$ for learning $\widetilde{\textrm{pr}}(A^{*}=1\mid Y=y,X=x)$, since the subset $\{(p^{(t)}_{i},x_{i}):Y_{i}=y\}$ might be too small for any specific value of $y$. Another aspect to consider is the choice of regression technique, where one has the liberty to select from a variety of time-tested machine learning methods. We generally advise keeping the hyperparameters of the selected method constant across iterations, rather than adjusting them dynamically. The rationale for this suggestion is that varying hyperparameters can significantly impede the convergence of the process. In our simulation and data application, we adopt kernel regression with T-learner in the M-step, except for $\widetilde{\textrm{pr}}(A^{*}=1\mid Y,X)$ with continuous $Y$, for which we use kernel regression with S-learner. See §7 for details.
    From the algorithm, we obtain two sets of values for $i=1,\ldots,n$: (1) $\widehat{f}(a^{*}\mid X_{i}),\widehat{E}(A\mid a^{*},X_{i}),\widehat{E}(Z\mid a^{*},X_{i}),\widehat{E}(Y\mid a^{*},X_{i})$; and (2) $\widehat{f}(1-a^{*}\mid X_{i}),\widehat{E}(A\mid 1-a^{*},X_{i}),\widehat{E}(Z\mid 1-a^{*},X_{i}),\widehat{E}(Y\mid 1-a^{*},X_{i})$. Plugging these estimated nuisance parameters into (17) yields $\widehat{\psi}_{a^{*}}$ and $\widehat{\psi}_{1-a^{*}}$. To determine whether $a^{*}$ is either equal to 00 or $1$—essentially deciding whether $\widehat{\psi}_{a^{*}}$ is equal to $\widehat{\psi}_{1}$ or $\widehat{\psi}_{0}$—we leverage Assumption 4. Recall that we have four alternative conditions in Assumption 4. We will focus on Condition 1 and Condition 2 in Assumption 4, both translating to $E(A\mid A^{*}=0,X)<E(A\mid A^{*}=1,X)$. Based on this, we conclude $\hat{a}^{*}=1$ if $\sum^{n}_{i=1}\widehat{E}(A\mid a^{*},X_{i})>\sum^{n}_{i=1}\widehat{E}(A\mid 1-a^{*},X_{i})$, and $\hat{a}^{*}=0$ otherwise.
    We now propose an estimator for $\psi_{1}=E(Y^{a^{*}=1})$ under Condition 1 or Condition 2 of Assumption 4. Let $d(X)=\arg\max_{j=1,2}E(A\mid A^{*}=a^{*}_{j},X)$, $S_{j}=\{A-E(A\mid 1-a^{*}_{j},X)\}\{E(A\mid a^{*}_{j},X)-E(A\mid 1-a^{*}_{j},X)\}^{-1}\{Z-E(Z\mid 1-a^{*}_{j},X)\}\{E(Z\mid a^{*}_{j},X)-E(Z\mid 1-a^{*}_{j},X)\}^{-1}\{f(a^{*}_{j}\mid X)\}^{-1}\{Y-E(Y\mid a^{*}_{j},X)\}+E(Y\mid a^{*}_{j},X)$, $\widehat{S}_{j}=\{A-\widehat{E}(A\mid 1-a^{*}_{j},X)\}\{\widehat{E}(A\mid a^{*}_{j},X)-\widehat{E}(A\mid 1-a^{*}_{j},X)\}^{-1}\{Z-\widehat{E}(Z\mid 1-a^{*}_{j},X)\}\{\widehat{E}(Z\mid a^{*}_{j},X)-\widehat{E}(Z\mid 1-a^{*}_{j},X)\}^{-1}\{\hat{f}(a^{*}_{j}\mid X)\}^{-1}\{Y-\widehat{E}(Y\mid a^{*}_{j},X)\}+\widehat{E}(Y\mid a^{*}_{j},X)$, we have that
    $\begin{split}\psi_{1}&=\sum^{2}_{j=1}E\Bigl{[}\mathbb{I}\{d(X)=j\}S_{j}\Bigr{]}=\mathbb{P}\left[\sum^{2}_{j=1}\mathbb{I}\{d(X)=j\}S_{j}\right].\end{split}$ (18)
    Plugging in the estimated nuisance functions from Algorithm 1 and $\hat{d}$ gives
    $\begin{split}\widehat{\psi}_{1}&=\sum^{2}_{j=1}\mathbb{P}_{n}\Bigl{[}\mathbb{I}\{\hat{d}(X)=j\}\widehat{S}_{j}\Bigr{]}=\mathbb{P}_{n}\left[\sum^{2}_{j=1}\mathbb{I}\{\hat{d}(X)=j\}\widehat{S}_{j}\right].\end{split}$ (19)
    To derive the asymptotic properties of $\widehat{\psi}_{1}$, we introduce the following margin condition.
    Assumption 5 (Margin condition).
    There exists $\epsilon_{1}>0$ such that $\epsilon_{1}<\textrm{pr}(A^{*}=1\mid X)<1-\epsilon_{1}$, $E(A\mid A^{*}=1,X)-E(A\mid A^{*}=0,X)>\epsilon_{1}$, $\left|E(Z\mid A^{*}=1,X)-E(Z\mid A^{*}=0,X)\right|>\epsilon_{1}$ almost surely.
    The first inequality is a common positivity condition in causal inference. The second and third inequalities require that $E(A\mid A^{*}=1,X)$ and $E(A\mid A^{*}=0,X)$, as well as $E(Z\mid A^{*}=1,X)$ and $E(Z\mid A^{*}=0,X)$, are well-separated.
    Theorem 5.
    Suppose that the nuisance functions are estimated from a separate independent sample, e.g., by cross-fitting. Moreover, assume Assumptions 2-5 hold, $Y$ is bounded, and
    1. 1.
      $\|\widehat{E}(Y\mid A^{*}=a^{*},X)-E(Y\mid A^{*}=a^{*},X)\|=o_{p}(1)$, $\|\widehat{E}(A\mid A^{*}=a^{*},X)-E(A\mid A^{*}=a^{*},X)\|=o_{p}(1)$, $\|\widehat{E}(Z\mid A^{*}=a^{*},X)-E(Z\mid A^{*}=a^{*},X)\|=o_{p}(1)$, $\|\widehat{f}(A^{*}=a^{*}\mid X)-f(A^{*}=a^{*}\mid X)\|=o_{p}(1)$ for $a^{*}=0,1$.
    2. 2.
      There exists $\epsilon_{2}>0$ such that $\min\{\widehat{E}(A\mid A^{*}=1,X)-\widehat{E}(A\mid A^{*}=0,X),\bigl{|}\widehat{E}(Z\mid A^{*}=1,X)-\widehat{E}(Z\mid A^{*}=0,X)\bigr{|},\hat{f}(A^{*}=1\mid X)\}>\epsilon_{2}$ almost surely.
    Then we have
    $\widehat{\psi}_{1}-\psi_{1}$ $=(\mathbb{P}_{n}-\mathbb{P})\left[\sum^{2}_{j=1}\mathbb{I}\{d(X)=j\}\phi_{a_{j}^{*}}\right]$
    $+\max_{a^{*}\in\{0,1\}}\Biggl{\{}O_{p}(\|\widehat{f}(a^{*}\mid X)-f(a^{*}\mid X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|)$
    $+O_{p}(\|\widehat{E}(A\mid a^{*},X)-E(A\mid a^{*},X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|)$
    $+O_{p}(\|\widehat{E}(Z\mid a^{*},X)-E(Z\mid a^{*},X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|)$
    $+O_{p}(\|\widehat{E}(A\mid 1-a^{*},X)-E(A\mid 1-a^{*},X)\|\|\widehat{E}(Z\mid 1-a^{*},X)-E(Z\mid 1-a^{*},X)\|)$
    $+O_{p}(\|\widehat{E}(A\mid a^{*}_{j},X)-E(A\mid a^{*}_{j},X)\|^{\alpha}_{\infty})\Biggr{\}}+o_{p}(n^{-1/2}),$
    where $\alpha$ is a positive integer.
[…]
    Remarkably, Theorem 5 does not require nuisance functions to converge at root-$n$ rates as it suffices for their pairwise product of convergence rates to be faster than root-$n$, and for the converge rate of $\widehat{E}(A\mid a^{*}_{j},X)$ raised to the $\alpha$ power to likewise be faster than root-$n$, where $\alpha$ is a user-specified positive integer. The choice of $\alpha$ appears to offer a cost-free advantage; however, there is indeed a trade-off between $\alpha$ and the margin constant $\epsilon_{1}$ in Assumption 5. If $\epsilon_{1}$ tends to zero, $\alpha$ must also tend to zero, otherwise the constant term $\epsilon_{1}^{-\alpha}$ scaling $\|\widehat{E}(A\mid a^{*}_{j},X)-E(A\mid a^{*}_{j},X)\|^{\alpha}_{\infty}$ will become unbounded (see the proof of Theorem 5 in the Supplementary Materials §LABEL:supp-sec:proof_asymptotic).
    6 Extention to other functionals
    In this section, we revisit the examples of §2 and derive corresponding influence functions in the hidden treatment model. Specifically, we continue to assume that the analyst has access to a surrogate $A$ and a proxy $Z$ for the hidden treatment $A^{*}$, that satisfy Assumptions 2, 3, and 4. We establish that one can readily leverage the same techniques as in §3-4 for identification, and we generically denote as $\psi^{*}$ the corresponding observed data functional for which we derive influence functions. The estimation of nuisance functions then follows in an analogous manner as above, details of which we omit for brevity. To present our results in a unified fashion, we note that all examples given in §2 have full data influence functions of the bi-linear form
    $\begin{split}\phi_{\mathrm{g}}^{F}(Y,A^{*},X;&\psi^{*})=\sum_{a^{*}=0,1}IF_{a^{*};\psi^{*}}=\sum_{a^{*}=0,1}\Bigl{\{}q_{0,a^{*}}(X)h_{0,a^{*}}(X)g_{1,a^{*}}(Y,X)\mathbb{I}(A^{*}=a^{*})\\ &+q_{0,a^{*}}(X)g_{2,a^{*}}(Y,X)\mathbb{I}(A^{*}=a^{*})+h_{0,a^{*}}(X)g_{3,a^{*}}(Y,X)+g_{4,a^{*}}(Y,X)\Bigr{\}},\end{split}$ (20)
    where $q_{0,a^{*}}(X)$ and $h_{0,a^{*}}(X)$ are unknown nuisance functions, $g_{j,a^{*}}$ are known functions of $(Y,X)$, and $IF_{a^{*};\psi^{*}},a^{*}=0,1$, belongs to the doubly robust class of influence functions studied in Robins et al., (2008). To simplify the exposition, we suppress the influence function dependence on $\psi^{*}$ throughout as $E[IF_{a^{*};\psi^{*}}]=0,a^{*}=0,1$. The next theorem characterizes for each influence function in the class above, a corresponding observed data influence function in the hidden treatment model. While we state the result for the class given above, the proof is more general as it considers an even larger class of doubly robust influence functions introduced in Ghassami et al., (2022), an important generalization which includes influence functions for causal functionals identified using the proximal causal inference framework (Miao et al.,, 2016, 2018; Tchetgen Tchetgen et al.,, 2024) that do not strictly fall within the above doubly robust class (Cui et al.,, 2024). The proof is deferred to the Supplementary Materials §LABEL:supp-sec:proof_extension.
    Theorem 6.
    Suppose that the influence functions for the parameter of interest $\psi^{*}$ in the full data model $\mathcal{M}^{*}$ are of the form in (20), indexed by functions $\mathrm{g}=(g_{1,a^{*}},g_{2,a^{*}},g_{3,a^{*}},g_{4,a^{*}},a^{*}=0,1)\in\mathcal{G}$ for a certain space of functions $\mathcal{G}$ in $L_{0}^{2}(X,Y)$ implied by the semiparametric model; then
    $\begin{split}\phi_{\mathrm{g}}&(Y,A,Z,X;\psi^{*})=\sum_{a^{*}=0,1}\Bigl{\{}q_{0,a^{*}}(X)h_{0,a^{*}}(X)g_{1,a^{*}}\left(Y,X\right)M_{a^{*},A}M_{a^{*},Z}\\ &+q_{0,a^{*}}(X)g_{2,a^{*}}\left(Y,X\right)M_{a^{*},A}M_{a^{*},Z}+h_{0,a^{*}}(X)g_{3,a^{*}}(Y,X)+g_{4,a^{*}}(Y,X)\Bigr{\}}\end{split}$ (21)
    is an influence function for $\psi^{*}$ in the hidden treatment model $\mathcal{M}$, where
    $\begin{split}M_{a^{*},A}=\{A-E(A\mid A^{*}=1-a^{*},X)\}\{E(A\mid A^{*}=a^{*},X)-E(A\mid A^{*}=1-a^{*},X)\}^{-1},\\ M_{a^{*},Z}=\{Z-E(Z\mid A^{*}=1-a^{*},X)\}\{E(Z\mid A^{*}=a^{*},X)-E(Z\mid A^{*}=1-a^{*},X)\}^{-1}.\end{split}$ (22)
    The set of influence functions in $\mathcal{M}$ is $\{\phi_{\mathrm{g}}:\mathrm{g}\in\mathcal{G}\}+\Lambda^{\bot}$, where $\Lambda^{\bot}$ is given in (13).
    Upon deriving the set of influence functions, an efficient estimator for $\psi^{*}$ in the hidden treatment model can be constructed as follows:
    1. 1.
      Obtain a preliminary (potentially inefficient) estimator $\hat{\psi}$ that solves $0=\mathbb{P}_{n}\hat{\phi}_{\mathrm{g}}(Y,A,Z,X)$ by using an estimated influence function $\hat{\phi}_{\mathrm{g}}(Y,A,Z,X)$ as an estimating function for a fixed choice $\mathrm{g}$.
    2. 2.
      Perform a one-step update in the direction of estimated EIF $\hat{\psi}^{eff}=\hat{\psi}+\mathbb{P}_{n}\hat{\phi}_{eff}(Y,A,Z,X),$ where $\hat{\phi}_{eff}(Y,A,Z,X)$ is an estimator of the EIF of ${\phi}_{eff}(Y,A,Z,X)=\Pi(\phi_{\mathrm{g}}\mid\Lambda)$ evaluated at $\hat{\psi}$ given in (LABEL:supp-eq:est_eif) in the Supplementary Materials §LABEL:supp-sec:closed_form_w.
    Example 2.
    ETT (continued). An observed data influence function is derived by replacing the unobserved indicator $\mathbb{I}(A^{*}=a^{*}),a^{*}=0,1$ in (2) with $M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)$ in (22), which we can write as
    $\phi_{ett}$ $=\frac{1}{\textrm{pr}(A^{*}=1)}M_{1,A}(A,X)M_{1,Z}(Z,X)\{Y-E(Y\mid A^{*}=1,X)\}$
    $\quad-\frac{1}{\textrm{pr}(A^{*}=1)}\frac{\textrm{pr}(A^{*}=1\mid X)}{1-\textrm{pr}(A^{*}=1\mid X)}M_{0,A}(A,X)M_{0,Z}(Z,X)\{Y-E(Y\mid A^{*}=0,X)\}$
    $\quad+\frac{E(Y\mid A^{*}=1,X)-E(Y\mid A^{*}=0,X)-\psi_{ett}}{\textrm{pr}(A^{*}=1)}\textrm{pr}(A^{*}=1\mid X).$
    Example 3.
    QTE (continued). The nuisance parameters $\eta_{1}(Y,X;\theta_{1})=\textrm{pr}(Y\leq\theta_{1}\mid X,A^{*}=a^{*})$ and $\eta_{2}(X)=\textrm{pr}(A^{*}=a^{*}\mid X)$ can both be identified by the approach in §3 by treating $\mathbb{I}(Y\leq\theta_{1}^{a^{*}})$ as a binary outcome.
    The influence function in the observed data model is obtained by substituting $\mathbb{I}(A^{*}=a^{*})$ in (4) with $M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)$ in (22), which gives
    $\phi_{a^{*},\gamma}=M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)\frac{\mathbb{I}(Y\leq\theta_{1}^{a^{*}})-\eta_{1}(X;\theta_{1}^{a^{*}})}{\eta_{2}(X)}+\eta_{1}(X;\theta_{1}^{a^{*}})-\gamma.$
    Example 4.
    Semi-linear structural regression model (continued). The full data influence function in (5) can be reformulated as $\phi^{F}_{sl}=\mathbb{I}(A^{*}=1)h(X)\{Y-\beta_{a^{*}}-\beta_{x}(X)\}\{1-\textrm{pr}(A^{*}=1\mid X)\}+\mathbb{I}(A^{*}=0)h(X)\{Y-\beta_{x}(X)\}\{-\textrm{pr}(A^{*}=1\mid X)\}.$ An influence function in the observed data model can be obtained by replacing $\mathbb{I}(A^{*}=a^{*}),a^{*}=0,1$ with $M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)$ in (22), which gives
    $\phi_{sl}$ $=M_{1,A}(A,X)M_{1,Z}(Z,X)h(X)\{Y-\beta_{a^{*}}-\beta_{x}(X)\}\{1-\textrm{pr}(A^{*}=1\mid X)\}$
    $\quad+M_{0,A}(A,X)M_{0,Z}(Z,X)h(X)\{Y-\beta_{x}(X)\}\{-\textrm{pr}(A^{*}=1\mid X)\}.$
    Example 5.
    Marginal structural generalized linear model (continued). The full data influence function in (5) can be rewritten as
    $\phi^{F}_{msm}$ $=\mathbb{I}(A^{*}=1)\frac{1}{\textrm{pr}(A^{*}=1\mid X)}h(1,V)\left\{Y-E(Y\mid A^{*}=1,X)\right\}$
    $\quad+\mathbb{I}(A^{*}=0)\frac{1}{\textrm{pr}(A^{*}=0\mid X)}h(0,V)\left\{Y-E(Y\mid A^{*}=0,X)\right\}$
    $\quad+\sum_{a^{*}}(-1)^{1-a^{*}}h(a^{*},V)\left\{E(Y\mid A^{*}=a^{*},X)-g^{-1}(\beta_{0}+\beta_{1}a^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\}.$
    An influence function in the observed data model can be derived by replacing $\mathbb{I}(A^{*}=a^{*}),a^{*}=0,1$ with $M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)$ in (22), which gives
    $\phi_{msm}$ $=\frac{M_{1,A}(A,X)M_{1,Z}(Z,X)}{\textrm{pr}(A^{*}=1\mid X)}h(1,V)\left\{Y-E(Y\mid A^{*}=1,X)\right\}$
    $\quad+\frac{M_{0,A}(A,X)M_{0,Z}(Z,X)}{\textrm{pr}(A^{*}=0\mid X)}h(0,V)\left\{Y-E(Y\mid A^{*}=0,X)\right\}$
    $\quad+\sum_{a^{*}}(-1)^{1-a^{*}}h(a^{*},V)\left\{E(Y\mid A^{*}=a^{*},X)-g^{-1}(\beta_{0}+\beta_{1}a^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\}.$
    7 Simulation study
    In this section, we evaluate the performance of our proposed estimator for both binary and continuous outcomes. The covariates $X=(X_{1},X_{2})$ are generated from a 2-dimensional uniform $[0,1]^{2}$ distribution. We then generate $A^{*}$ conditional on $X$ from a Bernoulli distribution with $\textrm{pr}(A^{*}=1\mid X)=1/\{1+\exp(-X_{1})+\exp(-X_{2})\}$. Next, we generate $A$ and $Z$ from $\textrm{pr}(A=1\mid A^{*}=1,X)=\text{logit}^{-1}(1+2X_{1})$, $\textrm{pr}(A=1\mid A^{*}=0,X)=\text{logit}^{-1}(-1+0.5X_{1}-X_{2})$, $\textrm{pr}(Z=1\mid A^{*}=1,X)=\text{logit}^{-1}(-1-X_{1}+0.5X_{2})$, $\textrm{pr}(Z=1\mid A^{*}=0,X)=\text{logit}^{-1}(2+X_{1})$, where $\text{logit}(p)=\ln p(1-p)^{-1}$.
    We consider two outcome regression models:
    i. (binary outcome) $\textrm{pr}(Y=1\mid A^{*}=1,X)=\text{logit}^{-1}(-2+0.5X_{1}+X_{2}),$
    $\textrm{pr}(Y=1\mid A^{*}=0,X)=\text{logit}^{-1}(1-2X_{1}+0.5X_{2});$
    ii. (continuous outcome) $Y=\sin(\pi X_{1})+(A^{*}-0.5)(X_{1}+X_{2})+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,1).$
    The specific form for our kernel regression model in M-step is given in the Supplementary Materials §LABEL:supp-sec:simu_details. Note that for $A$, we only need to estimate one of the probabilities $\textrm{pr}(A^{*}=1\mid A=a,X=x),a=0,1$, instead of both. This also applies to $\textrm{pr}(A^{*}=1\mid Z=z,X=x)$ for $z=0,1$ and $\textrm{pr}(A^{*}=1\mid Y=y,X=x)$ for $y=0,1$. In practice, one may elect to estimate the most prevalent surrogate subsample, i.e., we estimate $\textrm{pr}(A^{*}=1\mid A=1,X=x)$ if $n^{-1}\sum^{n}_{i=1}A_{i}\geq 0.5$, and $\textrm{pr}(A^{*}=1\mid A=0,X=x)$ if $n^{-1}\sum^{n}_{i=1}A_{i}<0.5$. We use 5-fold cross-fitting, and the number of Monte Carlo runs is 200. The threshold for convergence is set at 0.001.
    For comparison, we also present results for an infeasible estimator which substitutes the true values of nuisance parameters into (19), and the naive estimator, which treats $A$ as $A^{*}$ and fits a logistic regression of $Y$ on $(A,X)$ for binary $Y$, and a kernel regression for continuous $Y$. In our setting, $A$ matches $A^{*}$ around 82% of the time. The results are displayed in Table 1. From the results, one can see that our proposed estimator shows robust performance, achieving coverage close to the nominal rate of 95%, while the naive estimator is severely biased.
    bias(s.e.)(×$\times 100$) coverage
    $Y$ proposed infeasible naive proposed
    $n=1000$ $\psi_{1}$ $0.14(0.23)$ $-0.04(0.22)$ $8.49(0.15)$ 92.5%
    Binary $\psi_{0}$ $0.19(0.20)$ $0.13(0.19)$ $-4.15(0.15)$ 95%
    $n=2000$ $\psi_{1}$ $0.04(0.16)$ $-0.11(0.13)$ $8.51(0.10)$ 96%
    $\psi_{0}$ $-0.01(0.15)$ $-0.04(0.14)$ $-4.21(0.11)$ 95.5%
    $n=1000$ $\psi_{1}$ $2.55(1.55)$ $0.52(0.49)$ $-20.29(0.37)$ 97%
    Continuous $\psi_{0}$ $-1.03(1.70)$ $-0.36(0.45)$ $11.88(0.36)$ 93%
    $n=2000$ $\psi_{1}$ $1.68(0.38)$ $0.08(0.34)$ $-21.21(0.26)$ 96.5%
    $\psi_{0}$ $-1.03(0.38)$ $0.21(0.30)$ $11.91(0.25)$ 96%
    Table 1: Simulation results when $Y$ is binary or continuous. The “proposed” column is for our proposed estimator in (19). The “infeasible” column is for the estimator using true values of nuisance parameters. The “naive” column is for the estimator that treats $A$ as $A^{*}$. The coverage for the proposed estimator is given in the last column.
    8 Real data example
    Alzheimer’s Disease (AD) is a progressive neurodegenerative disorder that primarily affects the elderly, characterized by significant cognitive decline and memory impairment. One of the hallmark changes observed in AD is the atrophy of the hippocampus, a critical brain region involved in memory formation and spatial navigation. As the disease advances, neuronal damage occurs, resulting in brain atrophy that is especially pronounced in the hippocampus during the early stages of AD (Fox et al.,, 1996; Yu et al.,, 2022). We aim to analyze the impact of AD on hippocampal volume using data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI). The data usage acknowledgment is included in §LABEL:supp-sec:adni_acknowledgment of the supplement.
    Diagnosing AD involves a comprehensive approach, combining cognitive assessments such as the Mini-Mental State Examination (MMSE) and neuroimaging techniques like MRI and PET scans to detect brain changes. However, AD diagnosis may not reflect the actual latent disease status (Hunter et al.,, 2015). The diagnosis essentially serves as an approximation for the unobserved actual state of the disease.
    Our approach, as outlined in Section 5, requires selecting a surrogate and a proxy that are conditionally independent given a participant’s true but hidden AD status. Because AD diagnosis is essentially a composite of several measured factors including tests scores and biomarker measurements, using it as a surrogate would invalidate any such measurements as potential proxies given that they would fail to be independent. Instead, we use the recall and orientation subscores of the MMSE as indicators: one as a surrogate for underlying disease status and the other as a proxy. Specifically, the recall component evaluates the ability to remember three previously mentioned objects, while the orientation component assesses the awareness of the current time and place. We posit that the recall and orientation scores, which assess distinct cognitive functions, are independent given disease status and other covariates. This claim finds support in Yew et al., (2013), which revealed distinct neural correlates of AD: performance on orientation scores was associated with the posterior hippocampus, whereas memory performance was more affected by the anterior hippocampal regions. The scores are converted into binary variables, where 0 indicates a perfect score in a subcategory, and 1 indicates any deviation. In our analysis, we also account for important observed confounders such as age, sex, and number of years of education.
    We consider the underlying AD status at Month 6 as the latent exposure, and the outcome of interest is hippocampal volume at Month 12. The MMSE recall subscore at Month 6 serves as the surrogate, while the MMSE orientation subscore at the same time point acts as the proxy. We consider baseline age and sex as observed confounders. We include 1,172 subjects with complete MMSE scores at Month 6 and measurements of hippocampal volume at Month 12. Among these subjects, the hippocampal volume at Month 12 ranges from 2,839 to 11,041 mm3, with an average of 6,627 mm3 (SD = 1,258). Regarding the MMSE subscores, 50.4% of participants achieved a full score for orientation, and 45.3% achieved a full score for recall. The average baseline age is 73.4 years old (SD = 7.0); 44.4% of the subjects are female.
    We conduct separate analyses for the female and male groups, adjusting for age and years of education as potential confounders within each group. We consider the estimator proposed in equation (19) in §5. As in the simulation study, we use 5-fold cross-fitting. To address the influence of a particular random split on the estimates, we adopt the median adjustment method by repeating the cross-fitting 21 times (Chernozhukov et al.,, 2018).
    Our results indicate that for females, AD leads to a decrease of 1,756 mm3 in hippocampal volume, with 95% confidence interval of $[1,370,2,142]$. For males, AD leads to a decrease of 1,406 mm3 in hippocampal volume, with 95% confidence interval of $[853,1,960]$ (Fox et al.,, 1996). These findings appear to be aligned with prior research indicating that females are affected by AD more severely than males, and AD-control brain weight differences were significantly greater for females (Filon et al.,, 2016). For comparison, we also performed linear regression for both groups, with hippocampal volume as the outcome, error-prone diagnosis as the main exposure, age, and education years as regressors. These analyses indicate an AD-related decrease of 1,162 mm3 for females (95% CI: $[976,1,347]$) and 971 mm3 for males (95% CI: $[793,1,151]$), which are of smaller magnitude than our estimates, suggesting the potential for effect attenuation, a common consequence of exposure measurement error. In addition to linear regression, we also implemented the more complex doubly robust Targeted Maximum Likelihood Estimator (Benkeser et al.,, 2017), which is essentially biased in this context and yields results similar to those of linear regression.
    9 Discussion
    In this paper, we considered the fundamental problem of causal inference with a hidden binary treatment, an all too common problem in epidemiology and the social sciences. We proposed a unified solution which delivers nonparametric identification of a rich class of causal parameters, by drawing on a strategy which leverages a surrogate and a proxy variable of the hidden treatment to restore identification, while bypassing the need for validation data or restrictive functional form assumptions prominent in prior measurement error literature. An important contribution of the paper is to develop formal semiparametric theory which gives rise to influence function-based nonparametric estimators with important robustness and efficiency properties. The approach provides a generic roadmap for constructing novel semiparametric estimators in a large class of hidden treatment models. The paper also addresses the challenge of estimating high-dimensional nuisance functions defined in terms of the hidden treatment, therefore ruling out direct application of standard machine learning and nonparametric estimators. To resolve this difficulty, we proposed a semiparametric EM algorithm for estimating nuisance functions. We establish the semiparametric efficiency bound for discrete outcomes, and provide an approximation for continuous outcomes.
    The proposed methods may be improved or extended in several directions. First, our results can be extended to scenarios where the hidden treatment is a categorical variable with $k$ levels, which we anticipate will present several new challenges both for identification and semiparametric estimation. Additionally, extending these methods to continuous treatments presents a promising avenue, particularly given the practical relevance of continuous treatments in a wide range of scientific inquiries, such as for instance evaluating the impact of the magnitude of air pollution (e.g. levels of ambient PM2.5) on health outcomes.